A television antenna is situated atop a hill. From a point 200m from the base of the hill, the angle of elevation of the top of the antenna is 80º. The angle of elevation of the antenna from the same point is 75º. How tall is the antenna?
1
Expert's answer
2015-05-29T05:22:41-0400
The height of the antenna could be found by the following formula:
H = Dsin(80º) -D sin(75º) = D(sin(80º) - sin(75º) ), where D = 200 m
H = 200 (0.985 - 0.966) = 3.8 [m]